OpenStudy (anonymous):

write an equation of the line that passes through the given point and has the given slope m. (5,-5) m=-2 write an equation of the line that passes through the given point and has the given slope m. (5,-5) m=-2 @Mathematics

OpenStudy (anonymous):

y=mx+c, plug in x,y,m given to get the c then u get the equation required

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hey sis

OpenStudy (anonymous):

hellp

OpenStudy (anonymous):

y=mx+c is the general equation for linear line. given coord (x=5,y=-5) and slope,m=-2 u wil have -5=-2(5)+c after get the c, put in the c into equation y=-2x+c to get the answer

OpenStudy (anonymous):

oh ok so whats the answer

OpenStudy (anonymous):

amyy